OpenStudy (anonymous):

I will give you a medal!! Find an equation of the line that is tangent to the circle x^2 + y^2 = 25 at the point P(−4, −3). (Give your answer in slope intercept form.)

OpenStudy (aravindg):

did u try to take the derivative?

OpenStudy (cwrw238):

this is a circle whose centre is at (0,0) and the slope of the radius to thE poinT (-4,-3) IS -3/-4 = 3/4. the tangent at this point is perpendicular to the radius so its slope will be -4/3 use the geearal form y-y1 = m(x-x1) where m = the above slope and (x1,y1) = (-4,-3)